OpenStudy (anonymous):

Find the derivative of F(x)=3x^2 algebraically using the definition of the derivative.

OpenStudy (anonymous):

F(x)=3x^2 F(x) = 3(2)x^(2-1) F(x) = 6x^1 F(x) = 6x

OpenStudy (anonymous):

that is all that it is asking for when it said algebratically it confussed me

OpenStudy (anonymous):

are you currently in a section where you have to prove it using limits, epsilon proofs? the way i solved it above is how you normally solve it.

OpenStudy (anonymous):

we are put we just got done learning it the way that you just did it

OpenStudy (raden):

f ' = lim (h->0) ( f(x+h) - f(x) )/h

OpenStudy (anonymous):

I am so confussed

OpenStudy (anonymous):

Is the way that you did it algebrtically

OpenStudy (raden):

no, it is using the definition of the derivative. first, if given f(x) = 3x^2 then f(x+h) = ....

OpenStudy (anonymous):

right I need the algebratically using the definition of the derivative.

OpenStudy (raden):

got it ?

OpenStudy (anonymous):

I got f(3x^2+h)

OpenStudy (raden):

nope. if given f(x) = 3x^2 then f(x+h) = 3(x+h)^2 u need expand it : f(x+h) = 3(x+h)^2 = 3(x^2+2xh + h^2) = 3x^2 + 6xh + 3h^2

OpenStudy (anonymous):

ok

OpenStudy (raden):

now, apply of them to defined of derivative using limit above f '(x) = lim (h->0) ( f(x+h) - f(x) )/h = lim (h->0) (3x^2 + 6xh + 3h^2 - 3x^2 )/h = lim (h->0) (6xh+3h^2 )/h so far so good ?

OpenStudy (anonymous):

yes

OpenStudy (raden):

now, can u simplify that if all thing in numerator divided by h

OpenStudy (anonymous):

yes but you can;t simplify it anmore can you

OpenStudy (anonymous):

so wouldn't tht be the answer

OpenStudy (raden):

= lim (h->0) (6xh+3h^2 )/h = lim (h->0) 6xh/h + 3h^2/h = lim (h->0) 6x + 3h

OpenStudy (anonymous):

ok got it and that is the answer correct.

OpenStudy (raden):

just subtitute h=0, u got the answer right :)

OpenStudy (anonymous):

you mean so the answer is 6x+3(0)

OpenStudy (anonymous):

so then the answer is just 6x

OpenStudy (raden):

yup

OpenStudy (anonymous):

ok thanks

OpenStudy (raden):

that is how to get derivative by using definition, it is always using limit you're welcome

OpenStudy (anonymous):

thanks

OpenStudy (raden):

yw again :)